OpenStudy (anonymous):

I GIVE MEDAL PLZ HELP Question 1 (Multiple Choice Worth 1 points) [6.05] MA.7.G.4.4 A container states that it should be stored at a temperature below thirty-five degrees Celsius. Convert the temperature to degrees Fahrenheit. Round to the nearest whole number, if necessary. five degrees Fahrenheit thirty-seven degrees Fahrenheit ninety-five degrees Fahrenheit one hundred twenty-one degrees Fahrenheit < i put this but im not sure if its correct

OpenStudy (igreen):

35 Celsius is 95 Fahrenheit..you just have to convert it.

OpenStudy (anonymous):

so its c

OpenStudy (hauntedwoodsgal):

Lydia received a parking ticket because she parked ten feet from a fire hydrant. She had estimated her distance from the fire hydrant as being the minimum allowable distance, which is fifteen feet. What was her percent error? fifteen percent thirty-three percent fifty percent sixty-seven percent Question 18 (Multiple Choice Worth 1 points) [5.06] MA.7.A.1.2 Danika has saved enough money to pay cash for a used car. She wants to buy a car that costs five thousand five hundred dollars before taxes. If she will be charged six percent sales tax, what is the total amount she will pay for the car? five thousand one hundred seventy dollars five thousand eight hundred thirty dollars five thousand eight hundred fifty-one dollars six thousand one hundred dollars Question 19 (Multiple Choice Worth 1 points) [8.01] MA.7.P.7.2 Which of the following is an overlapping event with a probability of fifty percent? A spinner is divided into eight equal sections. Four are shaded red and four are shaded blue with each color alternating (red, then blue, then red…). The red sections are labeled B, F, M, and T. The blue sections are labeled A, E, O, and U. Spinning a vowel or blue. Spinning the letter A or the letter E. Spinning red or the letter U. Spinning the letter F or blue. Question 20 (Multiple Choice Worth 1 points) [7.02] MA.7.G.2.1 You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 5 inches and the radius is 6 inches. If none of the photos overlap, what is the area of photographs you need? Use three point one four for pi. ninety-four point two square inches one hundred eighty-eight point four square inches four hundred seventy-one square inches five hundred sixty-five point two square inches

OpenStudy (hauntedwoodsgal):

this is for example

OpenStudy (anonymous):

no need @Hauntedwoodsgal

OpenStudy (hauntedwoodsgal):

k

OpenStudy (anonymous):

i think its d im good with that thanks guys

OpenStudy (hauntedwoodsgal):

can I have a medal thought

OpenStudy (hauntedwoodsgal):

I have 0 medals :(

OpenStudy (igreen):

The formula is \(C = 5/9 (F-32)\) \(35 = 5/9F - 17.77)\) \(52.77 = 5/9F\) \(F \approx 95\)

OpenStudy (igreen):

D is incorrect, the correct answer is C.

OpenStudy (hauntedwoodsgal):

yeah I was looking at c

OpenStudy (anonymous):

but @iGreen gave me the answer help me with another then yes you may have a medal @Hauntedwoodsgal

OpenStudy (hauntedwoodsgal):

yay thank you

OpenStudy (igreen):

I gave you one. @Hauntedwoodsgal

OpenStudy (hauntedwoodsgal):

thank you igreen thank you 4u